OpenStudy (anonymous):

If the odds in favor of a particular event are 3 to 5, what is the probability that the event will occur?

OpenStudy (amistre64):

eight out comes, 3/8

OpenStudy (anonymous):

3/5 = 0.6

OpenStudy (anonymous):

3/8 is wrong. There are no 8 outcomes described in the problem.

OpenStudy (slaaibak):

Win:Lose 3:5 Win = 3/8 Lose= 5/8

OpenStudy (amistre64):

odds, measure success to failure ratio 3 out of 8 5 out of 8 3:5 odds

OpenStudy (anonymous):

I read the problem as the probability being 3 in 5. So, 3/5.

OpenStudy (amistre64):

yeah, thats a common error in odds to probabilities

OpenStudy (amistre64):

if I have 3 pennies and 5 dimes: the odds of picking a penny is 3:5 the probability of picking a penny is 3/8